OpenStudy (anonymous):

Jordan wants to solve the following system using the elimination method: 2x + 3y = 10 x + y = 7 What number should the equation x + y = 7 be multiplied by to eliminate y?

OpenStudy (anonymous):

@sugarplum15 I just tag you because your my homie ;-; But i know u suck at math too '-'

OpenStudy (anonymous):

opposites cancel whats the opposite of 2?

OpenStudy (anonymous):

2x + 3y = 10 ?(x + y = 7 )

OpenStudy (michele_laino):

hint: if I multiply the second equation by -3, I get: \[ \Large - 3x - 3y = - 21\]